# Condor: A Code Discriminator Integrating General Semantics with Code Details

### v1

@@ 6.2. Datasets for Code Discrimination
Current code discriminators commonly use code generation datasets as evaluation benchmarks. For example, CodeRanker, CodeT (Inala et al., 2022; Chen et al., 2023) have been evaluated on APPS, MBPP, HumanEval (i.e., code generation datasets for Python programming language) (Hendrycks et al., 2021; Austin et al., 2021; Chen et al., 2021b), and MBR-EXEC, Coder-Reviewer reranking (Shi et al., 2022; Zhang et al., 2023) have been evaluated on Spider (i.e., generation dataset for SQL code), NL2Bash (i.e., generation dataset for Bash code ) (Yu et al., 2018; Lin et al., 2018). Although these code-generation datasets can be used to evaluate the overall performance of discriminators, they fall short in assessing the discriminators’s true potential to detect subtle code details. In code generation datasets, inputs are typically natural language descriptions prompting the model to generate corresponding code. For the same functionality, there may be multiple implementation approaches, so the generated code is not always textually similar. Additionally, these datasets rely on outputs generated by LLMs, making their effectiveness dependent on the specific LLM used. Therefore, we construct the CodeNanoFix dataset from real user-submitted code, focusing on detailed changes in the code repair process to provide a more direct and effective evaluation of the discriminator’s performance in distinguishing code details.
6.3. Code Repair Datasets
@@
There are some studies focus on the construction of code repair datasets, such as Defects4J, Bugs.jar, Bugs2Fix, QuixBug, IntroClass, HumanEvalPack (Just et al., 2014; Saha et al., 2018; Lu et al., 2021; Lin et al., 2017; Le Goues et al., 2015; Muennighoff et al., 2023). However, these datasets are not designed to evaluate the discriminator’s ability to detect subtle code details. The reasons are as follows. (1) For repository-level repair datasets like Defects4J and Bugs.jar (Just et al., 2014; Saha et al., 2018), the process involves multiple stages, including defect localization, code repair, and patch detection. LLMs can hardly directly generate a series of code candidates to discriminate without specific fine-tuning or prompt design. In contrast, our CodeNanoFix dataset includes independent problem descriptions and code, allowing LLMs to directly generate fixed code for buggy code. Additionally, CodeNanoFix focuses on fine-grained code differences rather than extensive modifications, making it easier to evaluate the model’s potential. (2) The Bugs2Fix, QuixBug, IntroClass, and HumanEvalPack datasets are function-level repair datasets, but they are still not suitable for evaluating the discriminators. The Bugs2Fix (Lu et al., 2021) is a Java code repair dataset that includes erroneous Java code with variable anonymization. Due to the anonymization and lack of natural language descriptions of the code’s functionality, this dataset focuses more on syntactic and type errors rather than semantic errors. The QuixBug dataset (Lin et al., 2017) contains 40 erroneous Python code samples, with only a test set and no training set, aimed at evaluating code repair effectiveness. However, this dataset focuses on different repair problems and lacks fine-grained variations of code for the same repair task, making it challenging to evaluate the capabilities of discriminators. The IntroClass dataset (Le Goues et al., 2015) provides tens of thousands of erroneous C language answers for six programming tasks, while CodeNanoFix is sourced from thousands of different problems, offering greater semantic diversity. The HumanEvalPack (Muennighoff et al., 2023) dataset is based on 164 problems from the HumanEval (Chen et al., 2021b) dataset, with each problem containing a manually injected bug. However, these samples are created by intentionally introducing errors into correct code, rather than reflecting real-world bug-fixing scenarios, making it difficult to capture the true nature of code repair. Additionally, each problem in this dataset corresponds to only one bug and its corrected version, lacking coverage of diverse repair contexts for the same problem and thus failing to effectively showcase the variety of potential fixes. Therefore, we introduce a novel repair dataset, CodeNanoFix, that addresses the limitations of existing datasets in evaluating a model’s ability to recognize subtle code differences. In the CodeNanoFix dataset, the differences between the buggy and fixed code are minimal, enabling a more effective assessment of the model’s discriminative capability.
7. Conclusion
@@
8. Data Availability
Our code and dataset are available at the link: https://github.com/LIANGQINGYUAN/Condor

### v2

@@ VI-B Datasets for Code Discrimination
Current code discriminators commonly use code generation datasets as evaluation benchmarks. For example, CodeRanker, CodeT [coderanker, codet] have been evaluated on APPS, MBPP, HumanEval (i.e., code generation datasets for Python programming language) [apps, mbpp, humaneval], and MBR-EXEC, Coder-Reviewer reranking [mbrexec, coder_reviewer_reranking] have been evaluated on Spider (i.e., generation dataset for SQL code) and NL2Bash (i.e., generation dataset for Bash code ) [spider, nl2bash]. Although these datasets can be used to evaluate the overall performance of discriminators, they fall short in assessing the discriminators’s true potential to detect subtle code details. In code generation datasets, inputs are typically natural language descriptions prompting the model to generate corresponding code. For the same functionality, there may be multiple implementation approaches, so the generated code is not always textually similar. Additionally, these datasets rely on outputs generated by LLMs, making their effectiveness dependent on the specific LLM used. Therefore, we construct the CodeNanoFix dataset from real user-submitted code, focusing on detailed changes in the code repair process to provide a more direct and effective evaluation of the discriminator’s performance in distinguishing code details.
VI-C Code Repair Datasets
@@
There are some studies that focus on the construction of code repair datasets, such as Defects4J, Bugs.jar, Bugs2Fix, QuixBug, IntroClass, HumanEvalPack [defects4j, bugsjar, codexglue, quixbugs, introclass, commitpack_muennighoff2023octopack]. However, these datasets are not designed to evaluate the discriminator’s ability to detect subtle code details. The reasons are as follows. (1) For repository-level repair datasets like Defects4J and Bugs.jar [defects4j, bugsjar], the process involves multiple stages, including defect localization, code repair, and patch detection. LLMs can hardly directly generate a series of code candidates to discriminate without specific fine-tuning or prompt design. In contrast, our CodeNanoFix dataset includes independent problem descriptions and code, allowing LLMs to directly generate fixed code for buggy code. Additionally, CodeNanoFix focuses on fine-grained code differences rather than extensive modifications, making it easier to evaluate the model’s potential. (2) The Bugs2Fix, QuixBug, IntroClass, and HumanEvalPack datasets are function-level repair datasets, but they are still not suitable for evaluating the discriminators. The Bugs2Fix [codexglue] is a Java code repair dataset that includes erroneous Java code with variable anonymization. Due to the anonymization and lack of natural language descriptions of the code’s functionality, this dataset focuses more on syntactic and type errors rather than semantic errors. The QuixBug dataset [quixbugs] contains 40 erroneous Python code samples, with only a test set and no training set, aimed at evaluating code repair effectiveness. However, this dataset focuses on different repair problems and lacks fine-grained variations of code for the same repair task, making it challenging to evaluate the capabilities of discriminators. The IntroClass dataset [introclass] provides tens of thousands of erroneous C language answers for six programming tasks, while CodeNanoFix is sourced from thousands of different problems, offering greater semantic diversity. The HumanEvalPack [commitpack_muennighoff2023octopack] dataset is based on 164 problems from the HumanEval [humaneval] dataset, with each problem containing a manually injected bug. However, these samples are created by intentionally introducing errors into correct code, rather than reflecting real-world bug-fixing scenarios, making it difficult to capture the true nature of code repair. Additionally, each problem in this dataset corresponds to only one bug and its corrected version, lacking coverage of diverse repair contexts for the same problem and thus failing to effectively showcase the variety of potential fixes. Therefore, we introduce a novel repair dataset, CodeNanoFix, that addresses the limitations of existing datasets in evaluating a model’s ability to recognize subtle code differences. In the CodeNanoFix dataset, the differences between the buggy and fixed code are minimal, enabling a more effective assessment of the model’s discriminative capability.
VII Conclusion
@@
VIII Acknowledgment
This work is sponsored by the National Key Research and Development Program of China under Grant No. 2023YFB4503803, the National Natural Science Foundation of China under Grant No. 62232003, No. 62402482, and No. W2411051.